solve -3(5-y) + 2 to 4th power = 18y
Answer by checkley77(12569)

-3(5-y)+2^4=18y
-15+3y+16=18y
3y-18y=-16+15
-15y=-1
y=-1/-15
y=1/15 answer.
proof
-3(5-1/15)+16=18*1/15
-3(5*15-1)/15+16=18/15
-3(75-1)/15+16=18/15
-3*74/15+16=18/15
-222/15+16=18/15
-222/15+240/15=18/15
18/15=18/15